To be a lifeguard you must swim continuously for 500 yards doing either the front crawl or the breaststroke. You can swim the front crawl 45 yards per minute and the breaststroke 35 yards per minute. It takes you 12 minutes to swim 500 yards. How long did you spend swimming the front crawl?
Answer by stanbon(75887)

To be a lifeguard you must swim continuously for 500 yards doing either the front crawl or the breaststroke.
You can swim the front crawl 45 yards per minute and the breaststroke 35 yards per minute.
It takes you 12 minutes to swim 500 yards. How long did you spend swimming the front crawl?
-------------------
Front Crawl DATA:
distance = x yrds. ; rate = 45 yrd/min ; time = d/r = x/45 minutes
----------------
Breastroke DATA:
distance = 500-x yrds; rate = 35 yrs/min ; time = d/r = (500-x)/3 minutes
--------------------
EQUATION
tiime + time = 12 minutes:
x/45 + (500-x)/35 = 12
35x + 45(500-x) = 35*45*12
-10x + 45*500 = 35*45*12
-10x = -3600
x = 360 yds. (front crawl distance)
-----------
Front crawl time = 360 yds / (45 yds/mim) = 8 minutes
=================================
Cheers,
Stan H.